Five percent of truck drivers in Europe are women. Suppose 10 truck drivers are selected
randomly to be interviewed about quality of work conditions.
a) What is the probability that two of the drivers are women?
b) What is the probability that at least one will be a woman?
c) What is the probability that at most two of the drivers are women?
d) What is the probability that four of the drivers are men?

Answer by Boreal(15235)

2 are women 10C2*0.05^2*0.95^8=0.0746
at least 1 a woman is 1- 0 are women or 1-.95^10=0.4013
0 are women is 0.5987 prob.
1 is a woman is 10*0.05^1*0.05^9=0.3151
2 are women from above is 0.0746
those 3 possibilities have a combined probability of 0.9885
-
use a calculator with VARS2 scroll to binompdf(10,0.95,4) ENTER --use n p x. answer is 3 x 10^(-6), or about 0
also 10C4(0.95^4)(0.05^6)